# Adaptive Convolution for CNN-based Speech Enhancement Models

### v1

@@ Abstract
Deep learning-based speech enhancement methods have significantly improved speech quality and intelligibility. Convolutional neural networks (CNNs) have been proven to be essential components of many high-performance models. In this paper, we introduce adaptive convolution, an efficient and versatile convolutional module that enhances the model’s capability to adaptively represent speech signals. Adaptive convolution performs frame-wise causal dynamic convolution, generating time-varying kernels for each frame by assembling multiple parallel candidate kernels. A Lightweight attention mechanism leverages both current and historical information to assign adaptive weights to each candidate kernel, guiding their aggregation. This enables the convolution operation to adapt to frame-level speech spectral features, leading to more efficient extraction and reconstruction. Experimental results on various CNN-based models demonstrate that adaptive convolution significantly improves the performance with negligible increases in computational complexity, especially for lightweight models. Furthermore, we propose the adaptive convolutional recurrent network (AdaptCRN), an ultra-lightweight model that incorporates adaptive convolution and an efficient encoder-decoder design, achieving superior performance compared to models with similar or even higher computational costs.
Index Terms:
Speech enhancement, adaptive convolution, kernel attention, convolutional neural networks.
publicationid: pubid:
@@ I Introduction
Dynamic convolution is initially designed for CV tasks, where it aggregates information from an entire image to compute kernel attention weights, resulting in a fixed kernel configuration at the image level rather than the pixel level [21]. This design presents challenges when applied to SE tasks. Specifically, performing dynamic convolution on the spectrogram of an entire utterance introduces non-causality, making it unsuitable for real-time applications. Moreover, for highly non-stationary noisy speech, an utterance-level kernel configuration may be suboptimal for different signal segments or frames. One intuitive understanding is that the optimal convolution kernels for speech frames and noise frames are likely to differ significantly. Therefore, it is essential to develop causal dynamic convolution capable of adapting to the frame-level local features of the signal.
In this paper, we introduce adaptive convolution, a frame-wise causal dynamic convolution mechanism. Adaptive convolution generates time-varying kernels for each frame rather than relying on a fixed kernel configuration across different input features. Given that convolution kernels can be regarded as filters, the proposed method is analogous to adaptive filtering, as both approaches adjust filter coefficients in real time based on the statistical characteristics of the input signal. Note that the proposed method differs from the adaptive convolution commonly used in CV. In typical CV applications [25, 26, 27], adaptive convolution calculates the kernels at each spatial position across the entire feature map using modules with complex structures, which leads to a significant increase in computational overhead. In contrast, our low-complexity adaptive convolution performs frame-level kernel assembling, allowing for better utilization of the time-varying features of speech, where all frequency bins within an individual frame share the same kernel. Moreover, kernel generation of the adaptive convolution in CV often relies on external, task-specific information, whereas ours is directly guided by the input features.
@@
The adaptive convolution kernel is obtained by aggregating multiple parallel candidate kernels through frame-level attention weights derived from input features. This enables the convolution operation to adapt to frame-level speech spectral features, facilitating more efficient feature extraction and reconstruction. The attention mechanism plays a critical role in ensuring the effectiveness of adaptive convolution. We explore several channel modeling approaches for kernel attention, including more lightweight single-frame and multi-frame modeling, as well as higher-performance temporal modeling. Additionally, for convolutional blocks composed of multiple depthwise and pointwise convolution layers, we propose a multi-head mechanism to jointly generate kernel attention for these layers. This mechanism can also simultaneously generate optional input and output temporal channel attention maps, further enhancing the capability to model non-stationary speech signals.
@@
To validate its effectiveness, we conduct extensive experiments on several CNN-based models of diverse structures and scales, including DPCRN [10] of different sizes, DCCRN [9], GTCRN [7], and LiSenNet [16], by replacing vanilla convolutions with adaptive convolutions. Experimental results demonstrate that adaptive convolution is an efficient and versatile convolutional module that significantly enhances the performance of various models, particularly those with lower computational burdens. Although adaptive convolution increases the number of parameters, the resulting increment in computational complexity, which is more crucial for many applications, is negligible. Moreover, the results reveal that increasing the number of candidate convolution kernels can further improve model performance. Ablation experiments also highlight the advantages of the proposed adaptive convolution over non-causal, utterance-level dynamic convolution, underscoring the importance of frame-wise kernel adjustment. Additionally, by visualizing the kernel attention weights, we clearly demonstrate that adaptive convolution can effectively assign appropriate candidate kernels to frames with varying spectral features.
@@
To further illustrate the practical value of the proposed adaptive convolution, we propose an ultra-lightweight model, adaptive convolution recurrent network (AdaptCRN). Inspired by ConvNeXt [19] and StarNet [20] blocks, we design a lightweight and efficient encoder-decoder architecture that integrates adaptive convolution. We also incorporate strategies such as grouped RNN [7, 28] and spectral compression to further enhance performance and reduce model scale. Experimental results show that AdaptCRN achieves competitive performance with only 38M MACs per second and 135K parameters, outperforming several other SOTA lightweight models with similar or even significantly higher computational cost.
@@
This paper is organized as follows: Section II reviews the dynamic convolution. Section III introduces the proposed adaptive convolution, detailing its attention mechanism and implementation. Section IV presents the proposed ultra-lightweight model, AdaptCRN. Section V describes the experimental setup. Section VI reports the experimental results, verifying the effectiveness and generalization of adaptive convolution, evaluating the performance of AdaptCRN, and visualizing the relationship between the kernel attention of adaptive convolution and speech spectral features. Finally, Section VII draws the conclusions of the paper.
@@ II-A Dynamic Convolution
$W=\sum_{k=1}^{K}A_{k}W_{k},$ (1)
where $W$ represents the resulting dynamic kernel, $K$ is the number of candidate kernels, $W_{k},k=1,2,\cdots,K$ denotes the $k$-th candidate kernel consisting of $C_{\text{out}}$ filters $W_{k}^{m}\in\mathbb{R}^{K_{\text{h}}\times K_{\text{w}}\times C_{\text{in}}},m=1,2,\cdots,C_{\text{out}}$, and $A_{k}\in\mathbb{R}$ is the attention weight assigned to $W_{k}$. The attention weights are computed using a squeeze-and-excitation mechanism. Specifically, global spatial information is first compressed through global average pooling (GAP), followed by a fully connected (FC) layer, a rectified linear unit (ReLU) activation, another FC layer, and a softmax function. Due to the small size of the convolution kernels and the lightweight nature of the attention mechanism, DyConv is computationally efficient at the cost of an increased number of parameters. Dynamic convolution similarly handles bias, but for simplicity, this detail is not included in the formulas.
The softmax function ensures that the attention weights sum to one, as expressed by $\sum_{k=1}^{K}A_{k}=1$. This normalization helps constrain the kernel space, facilitating the learning of the attention model. During training, candidate kernels with low attention weights are often difficult to optimize. To address this, DyConv incorporates temperature annealing in the softmax function, encouraging a near-uniform attention distribution in the early training stages. Specifically, the features are scaled by a temperature coefficient, which varies during training, before applying the softmax function. A higher temperature produces a flatter distribution of attention weights, facilitating the learning of all kernels. As training progresses, the temperature is gradually reduced, allowing the model to assign more distinct attention weights and improving its ability to select the most relevant kernels.
@@ II-B Omni-dimensional Dynamic Convolution
ODConv [23] introduces a multi-dimensional attention mechanism to enhance the performance of DyConv. It employs a parallel strategy to learn complementary attention weights for convolutional kernels across four dimensions of the kernel space: spatial, channel (input channel), filter (output channel), and candidate kernel dimensions. Specifically, in addition to kernel attention, ODConv assigns distinct attention weights, denoted as $A_{k}^{\text{s}}\in\mathbb{R}^{K_{\text{h}}\times K_{\text{w}}}$, $A_{k}^{\text{c}}\in\mathbb{R}^{C_{\text{in}}}$, and $A_{k}^{\text{f}}\in\mathbb{R}^{C_{\text{out}}}$, to different spatial locations, input channels, and output channels of the $k$-th candidate kernel. It can be represented as
$W=\sum_{k=1}^{K}A_{k}\left(A_{k}^{\text{s}}*A_{k}^{\text{c}}*A_{k}^{\text{f}}*W_{k}\right),$ (2)
where * denotes element-wise multiplication across different dimensions. The attention weights for each dimension are computed using a multi-head mechanism. The additional spatial, channel, and filter attention weights are all generated using the sigmoid function. For simple implementation, these weights can be shared across all candidate kernels, allowing the subscript $k$ of the corresponding attention descriptor in Eq. 2 to be omitted. By leveraging this refined kernel adjustment mechanism, ODConv achieves comparable or superior performance to standard DyConv while requiring fewer convolutional kernels and significantly fewer extra parameters.
@@ III Adaptive Convolution
Performing dynamic convolution on the spectrogram of an entire utterance inevitably introduces non-causality, and the resulting kernels may fail to adapt to the significant differences between distinct signal segments or frames. Therefore, we propose adaptive convolution, a frame-wise causal dynamic convolution mechanism, which enables the convolution operation to adapt to frame-level speech spectral features. Similar to adaptive filtering, adaptive convolution dynamically adjusts the convolutional kernel in real time based on the statistical characteristics of the input signal. As depicted in Fig. 1(a), the adaptive kernel is generated by aggregating multiple parallel candidate kernels with frame-level attention weights derived from input features, represented as
$W\left(t\right)=\sum_{k=1}^{K}A_{k}\left(t\right)W_{k},$ (3)
where $t$ denotes the frame index, emphasizing the time-varying nature of the kernel. The adaptive kernel is then applied in a causal convolution with the features of the corresponding frame to produce the output. For simplicity, we keep the biases of the convolution layers static rather than adaptive.
@@ III-A Kernel Attention
The attention mechanism is pivotal to the effectiveness of adaptive convolution, as it assigns attention weights to each parallel candidate kernel, determining the kernel composition for the current frame. As shown in Fig. 1(b), kernel attention aggregates the input features $Y\in\mathbb{R}^{C\times T\times F}$ along the frequency dimension through power average pooling to derive the descriptor $P\in\mathbb{R}^{C\times T}$ for the time-channel energy distribution, where $C,T,F$ represent the number of channels, frames and frequency bins, respectively. The corresponding pooling is formulated as
@@
where $c$ and $f$ denote the channel and frequency bin indices, respectively. $P$ is subsequently processed through channel modeling and a softmax function to compute the time-varying attention weights $A_{k}(t)$ for each candidate kernel. During training, we omit the temperature annealing strategy described in Section II-A for the softmax function, as our experiments indicate that it does not improve kernel optimization in adaptive convolution. This is likely because different candidate kernels dominate different frames, and the attention weights for many frames do not exhibit one-hot characteristics, eusuring sufficient optimization of all kernels even without temperature annealing.
For causal convolution, an intuitive channel modeling approach is to apply squeeze-and-excitation mechanism to the feature of the current frame. The compressed feature $P$ is processed through an FC layer, a ReLU activation, and another FC layer sequentially. Notably, in the deeper layers of the model, the single-frame channel modeling approach implicitly leverages historical information. This is because preceding layers, such as convolution layers with large temporal receptive fields and inter-frame RNNs, may have already incorporated information from previous frames. Additionally, we explore a multi-frame channel modeling approach, which consists of a stack of a one-dimensional (1-D) convolution layer, a ReLU activation, and an FC layer. The 1-D convolution explicitly incorporates information from previous frames, thereby extending the receptive field of channel modeling. Inspired by TRA [7], a lightweight temporal channel attention mechanism, we adopt a similar strategy by using gated recurrent unit (GRU) to model channel features along the time dimension. Specifically, a GRU followed by an FC layer is employed for temporal channel modeling. This leads to better utilization of inter-frame correlations, thus yielding more efficient adaptive kernels, albeit at the cost of introducing more parameters compared to single-frame modeling.
Benefiting from the lightweight design of kernel attention, adaptive convolution introduces negligible computational complexity to the model, which is critical in practical applications. However, this comes with an increase in parameters, as each adaptive convolution layer contains $K$ parallel candidate convolution kernels, resulting in a convolutional parameter count that is $K$ times the original.

### v2

@@ Abstract
Deep learning-based speech enhancement methods have significantly improved speech quality and intelligibility. Convolutional neural networks (CNNs) have been proven to be essential components of many high-performance models. In this paper, we introduce adaptive convolution, an efficient and versatile convolutional module that enhances the model’s capability to adaptively represent speech signals. Adaptive convolution performs frame-wise causal dynamic convolution, generating time-varying kernels for each frame by assembling multiple parallel candidate kernels. A lightweight attention mechanism is proposed for adaptive convolution, leveraging both current and historical information to assign adaptive weights to each candidate kernel. This enables the convolution operation to adapt to frame-level speech spectral features, leading to more efficient extraction and reconstruction. We integrate adaptive convolution into various CNN-based models, highlighting its generalizability. Experimental results demonstrate that adaptive convolution significantly improves the performance with negligible increases in computational complexity, especially for lightweight models. Moreover, we present an intuitive analysis revealing a strong correlation between kernel selection and signal characteristics. Furthermore, we propose the adaptive convolutional recurrent network (AdaptCRN), an ultra-lightweight model that incorporates adaptive convolution and an efficient encoder-decoder design, achieving superior performance compared to models with similar or even higher computational costs.
publicationid: pubid: 2998-4173 © 2025 IEEE. All rights reserved, including rights for text and data mining, and training of artificial intelligence and similar technologies. Personal use is permitted, but republication/redistribution requires IEEE permission. See https://www.ieee.org/publications/rights/index.html for more information.
@@ I Introduction
Dynamic convolution is initially designed for CV tasks, where it aggregates information from an entire image to compute kernel attention weights and applies a fixed kernel configuration at the image level rather than the pixel level [21]. A related approach, adaptive convolution, has also been proposed in the CV field, generating pixel-level independent kernels for each spatial location in the feature map [25, 26, 27]. However, these methods are not directly suitable for causal SE. The former introduces strong non-causality, while the latter typically relies on complex architectures or additional information for kernel computation, leading to excessive computational cost. Moreover, for highly non-stationary noisy speech, an utterance-level kernel configuration may be suboptimal for different signal segments or frames. One intuitive understanding is that the optimal convolution kernels for speech frames and noise frames are likely to differ significantly. This suggests that, for real-time SE, frame-level causal kernel adjustment could better adapt to time-varying signal characteristics while maintaining low computational overhead. This idea has been briefly explored in dynamic gate convolutional networks (DGCN) [28], where the global pooling in the squeeze-and-excitation mechanism of dynamic convolution is replaced with pooling along the frequency dimension within each frame to produce time-varying kernels. However, in DGCN, this mechanism is applied only to a subset of convolutional modules, without further refinement, systematic exploration, or extensive experimentation. In contrast to DGCN, our work thoroughly explores, refines, and extends the attention mechanism of this approach, demonstrating its generalizability across diverse baseline models. In particular, we emphasize its critical role in improving the performance of lightweight networks and simultaneously propose an ultra-lightweight model for real-time SE. Furthermore, we reveal a key connection between dynamic kernel adjustment and speech spectral characteristics. The main contributions of this paper are summarized in the following paragraphs.
@@
In this paper, we propose adaptive convolution, an efficient and versatile convolution mechanism that performs frame-wise causal dynamic convolution through an enhanced attention mechanism explicitly leveraging historical information. Adaptive convolution generates time-varying kernels for each frame rather than relying on a fixed kernel configuration across different input features. Given that convolution kernels can be regarded as filters, the proposed method is analogous to adaptive filtering, as both approaches adjust filter coefficients in real time based on the statistical characteristics of the input signal. The adaptive convolution kernel is obtained by aggregating multiple parallel candidate kernels through frame-level attention weights derived from input features. This enables the convolution operation to adapt to frame-level speech spectral features, facilitating more efficient feature extraction and reconstruction. The attention mechanism plays a critical role in ensuring the effectiveness of adaptive convolution. We explore several channel modeling approaches for kernel attention, including more lightweight single-frame and multi-frame modeling, as well as higher-performance temporal modeling. Additionally, for convolutional blocks composed of multiple depthwise and pointwise convolution layers, we propose a multi-head mechanism to jointly generate kernel attention for these layers. This mechanism can also simultaneously generate optional input and output temporal channel attention maps, further enhancing the capability to model non-stationary speech signals.
@@
In this paper, we highlight the effectiveness and generalizability of adaptive convolution across diverse model architectures. We conduct extensive experiments on several CNN-based models of diverse structures and scales, including DPCRN [10] of different sizes, DCCRN [9], GTCRN [7], and LiSenNet [16], by replacing vanilla convolutions with adaptive convolutions. Experimental results demonstrate that adaptive convolution is an efficient and versatile convolutional module that significantly enhances the performance of various models, particularly those with lower computational burdens. Although adaptive convolution increases the number of parameters, the resulting increment in computational complexity, which is more crucial for many applications, is limited. Ablation experiments also highlight the advantages of the proposed adaptive convolution over non-causal, utterance-level dynamic convolution, underscoring the importance of frame-wise kernel adjustment. Additionally, through a detailed analysis of the kernel attention weights, we clearly demonstrate that adaptive convolution can effectively assign appropriate candidate kernels to frames with varying spectral features.
@@
To further illustrate the practical value of the proposed adaptive convolution, we propose an ultra-lightweight model, adaptive convolution recurrent network (AdaptCRN). Inspired by ConvNeXt [19] and StarNet [20] blocks, we design a lightweight and efficient encoder-decoder architecture that integrates adaptive convolution. We also incorporate strategies such as grouped RNN [7, 29] and spectral compression to further enhance performance and reduce model scale. Experimental results show that AdaptCRN achieves competitive performance with only 41M MACs per second and 135K parameters, outperforming several other SOTA lightweight models with similar or even significantly higher computational cost.
@@
This paper is organized as follows: Section II reviews the dynamic convolution. Section III introduces the proposed adaptive convolution, detailing its attention mechanism and implementation. Section IV presents the proposed ultra-lightweight model, AdaptCRN. Section V describes the experimental setup. Section VI reports the experimental results, verifying the effectiveness and generalization of adaptive convolution, evaluating the performance of AdaptCRN, and demonstrating the relationship between the kernel attention of adaptive convolution and speech spectral features. Finally, Section VII draws the conclusions of the paper.
@@ II-A Dynamic Convolution
$\mathbf{W}=\sum_{k=1}^{K}A_{k}\mathbf{W}_{k},$ (1)
where $\mathbf{W}$ represents the resulting dynamic kernel, $K$ is the number of candidate kernels, $\mathbf{W}_{k},k=1,2,\cdots,K$ denotes the $k$-th candidate kernel consisting of $C_{\text{out}}$ filters $\mathbf{W}_{k}^{m}\in\mathbb{R}^{K_{\text{h}}\times K_{\text{w}}\times C_{\text{in}}},m=1,2,\cdots,C_{\text{out}}$, and $A_{k}\in\mathbb{R}$ is the attention weight assigned to $\mathbf{W}_{k}$. The attention weights are computed using a squeeze-and-excitation mechanism. Specifically, global spatial information is first compressed through global average pooling (GAP), followed by a fully connected (FC) layer, a rectified linear unit (ReLU) activation, another FC layer, and a softmax function. Due to the small size of the convolution kernels and the lightweight nature of the attention mechanism, DyConv is computationally efficient at the cost of an increased number of parameters. Dynamic convolution similarly handles bias, but for simplicity, this detail is not included in the formulas.
The softmax function ensures that the attention weights sum to one, as expressed by $\sum_{k=1}^{K}A_{k}=1$. As pointed out in DyConv, this normalization helps constrain the kernel space, facilitating the learning of the attention model. During training, candidate kernels with low attention weights are often difficult to optimize. To address this, DyConv incorporates temperature annealing in the softmax function, encouraging a near-uniform attention distribution in the early training stages. Specifically, the features are scaled by a temperature coefficient, which varies during training, before applying the softmax function. A higher temperature produces a flatter distribution of attention weights, facilitating the learning of all kernels. As training progresses, the temperature is gradually reduced, allowing the model to assign more distinct attention weights and improving its ability to select the most relevant kernels.
@@ II-B Omni-dimensional Dynamic Convolution
ODConv [23] introduces a multi-dimensional attention mechanism to enhance the performance of DyConv. It employs a parallel strategy to learn complementary attention weights for convolutional kernels across four dimensions of the kernel space: spatial, channel (input channel), filter (output channel), and candidate kernel dimensions. Specifically, in addition to kernel attention, ODConv assigns distinct attention weights, denoted as $\mathbf{A}_{k}^{\text{s}}\in\mathbb{R}^{K_{\text{h}}\times K_{\text{w}}}$, $\mathbf{A}_{k}^{\text{c}}\in\mathbb{R}^{C_{\text{in}}}$, and $\mathbf{A}_{k}^{\text{f}}\in\mathbb{R}^{C_{\text{out}}}$, to different spatial locations, input channels, and output channels of the $k$-th candidate kernel. It can be represented as
$\mathbf{W}=\sum_{k=1}^{K}A_{k}\left(\mathbf{A}_{k}^{\text{s}}*\mathbf{A}_{k}^{\text{c}}*\mathbf{A}_{k}^{\text{f}}*\mathbf{W}_{k}\right),$ (2)
where * denotes element-wise multiplication, with implicit broadcasting of the attention weights to align with the shape of $\mathbf{W}_{k}$. The attention weights for each dimension are computed using a multi-head mechanism. Specifically, the final FC layer in the squeeze-and-excitation attention module described in Section II-A is extended to 4 parallel FC layers, with the additional 3 layers followed by sigmoid functions to generate the spatial, channel, and filter attention weights. For simple implementation, these weights can be shared across all candidate kernels, allowing the subscript $k$ of the corresponding attention descriptor in Eq. 2 to be omitted. By leveraging this refined kernel adjustment mechanism, ODConv achieves comparable or superior performance to standard DyConv while requiring fewer convolutional kernels and significantly fewer extra parameters.
@@ III Adaptive Convolution
Considering the time-varying nature of speech signals and the real-time SE requirements for causality and computational efficiency, we propose adaptive convolution, a frame-wise causal dynamic convolution mechanism, which enables the convolution operation to adapt to frame-level speech spectral features. Similar to adaptive filtering, adaptive convolution dynamically adjusts the convolutional kernel in real time based on the statistical characteristics of the input signal. As depicted in Fig. 1(a), the adaptive kernel is generated by aggregating multiple parallel candidate kernels with frame-level attention weights derived from input features, represented as
$\mathbf{W}\left(t\right)=\sum_{k=1}^{K}A_{k}\left(t\right)\mathbf{W}_{k},$ (3)
where $t$ denotes the frame index, emphasizing the time-varying nature of the kernel. The learnable candidate kernels, $\mathbf{W}_{k},k=1,2,\cdots,K$, are end-to-end optimized and then fixed after training. The adaptive kernel is then applied in a causal convolution with the features of the corresponding frame to produce the output. For simplicity, we keep the biases of the convolution layers static rather than adaptive.
@@ III-A Kernel Attention
The attention mechanism is pivotal to the effectiveness of adaptive convolution, as it assigns attention weights to each parallel candidate kernel, determining the kernel composition for the current frame. As shown in Fig. 1(b), kernel attention aggregates the input features $\mathbf{Y}\in\mathbb{R}^{C\times T\times F}$ along the frequency dimension through power average pooling to derive the descriptor $\mathbf{P}\in\mathbb{R}^{C\times T}$ for the time-channel energy distribution, where $C,T,F$ represent the number of channels, frames and frequency bins, respectively. The corresponding pooling is formulated as
@@
where $c$ and $f$ denote the channel and frequency bin indices, respectively. $\mathbf{P}$ is subsequently processed through channel modeling and a softmax function to compute the time-varying attention weights $A_{k}(t)$ for each candidate kernel.
For causal convolution, an intuitive channel modeling approach is the single-frame channel modeling strategy. The compressed feature $\mathbf{P}$ is processed through an FC layer, a ReLU activation, and another FC layer sequentially, which essentially corresponds to the standard squeeze-and-excitation mechanism. Notably, the single-frame channel modeling approach may implicitly leverage historical information when adaptive convolution is used as a layer within the overall model. This is because the input features at each frame may already encode information from previous frames, depending on the structure of the preceding layers. Additionally, we explore a multi-frame channel modeling approach, which consists of a stack of a one-dimensional (1-D) convolution layer, a ReLU activation, and an FC layer. The 1-D convolution explicitly incorporates information from previous frames, thereby extending the receptive field of channel modeling. Inspired by TRA [7], a lightweight temporal channel attention mechanism, we adopt a similar strategy by using gated recurrent unit (GRU) to model channel features along the time dimension. Specifically, a GRU followed by an FC layer is employed for temporal channel modeling. This leads to better utilization of inter-frame correlations, thus yielding more efficient adaptive kernels, albeit at the cost of introducing more parameters compared to single-frame modeling.
@@
Benefiting from the lightweight design of kernel attention, adaptive convolution introduces limited computational complexity to the model when the original convolution kernel scale is not particularly large, which is critical in practical applications. However, this comes with an increase in parameters, as each adaptive convolution layer contains $K$ parallel candidate convolution kernels, resulting in a convolutional parameter count that is $K$ times the original. Moreover, the additional computational load primarily arises from the kernel aggregation step. When the original model has large convolution kernels, the resulting computational burden can become considerable, imposing certain constraints on its applicability.